What is 5 feet wide by 63 feet long by 3 inches deep in liquid volume. We are planning to lay a sidewalk and are having a little difficulty trying to sum up how much cement to buy. Thank you for your time. Raquel
Answer by stanbon(75887)

What is 5 feet wide by 63 feet long by 3 inches deep in liquid volume. We are planning to lay a sidewalk and are having a little difficulty trying to sum up how much cement to buy.
---
You buy cement is cubic yards.
---
Your Problem:
5 ft = (5/3) yds
---
63 ft = (63/3) = 21 yds
----
3 inches = (1/4)ft = (1/12) yds
-----
Total volume = (5/3)(21)(1/12) = 105/36 = 2.92 cubic yds
============================
Get at least 3 cubic yds.
============================
Cheers,
Stan H.